# Meta-Transfer Derm-Diagnosis: Exploring Few-Shot Learning and Transfer Learning for Skin Disease Classification in Long-Tail Distribution

### v1

@@ 1 Shot 5 Shot 1 Shot 5 Shot
SCAN Conv6 77.64±1.50 88.28±1.03 54.07±1.24 74.73±0.92
NCA WRN-28-10 71.27±1.50 84.23±1.19 45.91±1.08 62.83±1.01
Baseline 76.64±1.56 89.66±0.97 52.54±1.11 74.71±0.96
S2M2_R 77.15±1.59 90.97±0.89 55.49±1.13 78.17±0.84
NegMargin 77.98±1.45 90.65±0.92 56.04±1.14 77.75±0.87
PT+NCM 78.86±1.47 90.90±0.93 56.91±1.11 78.12±0.88
PEMbE-NCM 78.70±1.49 90.94±0.95 57.42±1.11 78.78±0.90
EASY 79.44±1.51 91.43±0.96 57.77±1.12 79.53±0.89
SCAN 81.21±1.46 92.08±0.85 58.75±1.14 81.43±0.77
DTL-Base(Ours) MobileNetV2 80.54±0.53 94.10±0.29 61.32±0.40 86.04±0.26
DTL-CutMix(Ours) 81.10±0.53 94.51±0.28 62.78±0.41 87.05±0.26
DTL-MixUp(Ours) 80.34±0.53 94.20±0.28 61.71±0.41 86.61±0.26
DTL-All-Augment(Ours) 81.44±0.53 94.76±0.28 63.22±0.41 87.48±0.25
DTL-ResizeMix(Ours) 83.06±0.51 95.05±0.27 65.40±0.40 88.08±0.25
Note:(1) The results of the SOTA models are taken from the SCAN [19]. (2) The complete references for the mentioned works in the table are as follows: PCN [14], SCAN [19], NCA [53], Baseline[47], S2M2_R [46], NegMargin [54], PT+NCM [48], PEM bE-NCM [53], EASY [55].
The SD-198 dataset contains 198 classes and consists only of clinical images. In contrast, the Derm7pt dataset has 18 classes with a mix of clinical and dermoscopic images. The ISIC2018 dataset, with its 7 classes, mainly features dermoscopic images. Although each of these three datasets shows a long-tail distribution, their unique features affect how models perform. Additionally, the number of parameters and the performances of the backbone models also play essential role in the overall model performance. For instance, MobileNetV2 has approximately 3.4 million parameters, while DenseNet121 boasts 8.1 million parameters. Furthermore, while episodic learning is employed in FEL and FETL models, transfer learning based training is utilized in DTL and DL models, a distinction that significantly impacts model performance. As can be observed in Figure 5, in all three datasets, FETL models consistently outperform FEL models, while DTL models demonstrate superior performance compared to DL models. Except for some rare cases, depending on the number of shots and dataset conditions, our experimental results indicate that the most successful approach in the proposed framework is the deep transfer learning based DTL model. As the number of shots increases, the performance of transfer learning-based DTL and DL models improves.
@@
When comparing our models without utilizing augmentation techniques, datasets containing relatively large number of classes, such as SD-198 and Derm7pt, exhibit similar behaviors. Episodic learning-based models tend to be more successful when the number of shots is low (i.e. 1 or 2 shots), while DTL and DL models show a tendency to outperform as the number of shots increases. For instance, with MobilNetV2 model on the SD-198 dataset, for a 5-Way 1-Shot scenario, the FETL model achieved 65.85%, whereas the DTL model achieved 64.06%. For a 5-Way 10-Shot scenario, the FETL model achieved 86.36%, whereas the DTL model achieved 91.37% (Table 1). The behaviour of the models are similar in Derm7pt dataset as well (Table 2). The reason behind this phenomenon lies in the adaptability of episodic learning, which performs better in scenarios with fewer shots due to the large number of classes. Additionally, clinical images inherently encapsulate various differences, making models trained in an episodic manner more inclined to tolerate these diversities. As the number of shots increases, models trained non-episodically become more successful in classifications.
@@
Since the ISIC2018 dataset comprises dermoscopic data, the discriminative power of domain-specific features becomes more crucial. Hence, transfer learning-based DTL and DL models outperform FETL and FEL models. For instance, for MobileNetV2, the accuracy rates are 58.49% for 2-Way 1-Shot with FETL and 64.83% with DTL, and 68.82% for 2-Way 10-Shot with FETL and 85.18% with DTL (Table 3). Episodic learning-based methods tend to remain superficial in training. Additionally, in parallel with the SD-198 dataset, models tend to become more successful as the number of shots increases in ISIC2018 and Derm7p datasets. The MobileNetV2 and DenseNet121 models tend to yield complementary outcomes. DenseNet121, with its greater parameter count, is more susceptible to overfitting while training. Conversely, MobileNetV2, with its simpler architecture, is better suited for practical applications in this setting.
@@ 5.3 Comparison with Current State-of-the-Art
Table 5: Comparison of our models and the SOTA methods. Values in the table are Accuracies (%) of the corresponding models on the ISIC2018 dataset.
Method Backbone 2 Way 3 Way
3 Shot 5 Shot 10 Shot 3 Shot 5 Shot 10 Shot
Meta-DermDiagnosis Conv6 64.50 73.50 79,70 - - -
MetaMed - Transf. Learn. Conv4 66.88 73.88 81.37 54.83 59.33 69.75
MetaMed - Normal Aug. 72.75 75.62 81.37 54.83 59.33 69.75
MetaMed - CutOut 70.37 77.62 81.87 55.5 65.41 69.75
MetaMed - MixUp 75.37 78.25 84.25 58.5 61.25 71
MetaMed - CutMix 73.25 76.87 80.62 58.66 61.5 66.5
PT-MAP WRN 68.15 70.87 74.19 53.17 55.61 59.57
Baseline+ 64.77 70.27 74.67 53.2 54.16 57.87
NegMargin 71.33 72.67 75.17 60.69 57.58 63.04
Baseline 68.77 71.03 76.97 56.8 59.2 65.22
PFEMed 81.69 83.87 85.14 66.94 69.78 73.81
DTL-Base(Ours) MobileNetV2 77.26 81.63 85.18 63.98 69.35 75.01
DTL-CutMix(Ours) 77.71 81.79 85.97 64.37 69.86 75.73
DTL-MixUp(Ours) 79.02 82.95 86.4 66.84 71.15 76.48
DTL-ResizeMix(Ours) 78,02 82.75 86.69 65.56 70.87 76.97
DTL-All-Augment(Ours) 78,96 83.21 86.83 66.12 71.28 76.94
Note: (1) The results of the SOTA models are taken from the PFEMed [25]. (2) The complete references for the mentioned works in the table are as follows: Meta-DermDiagnosis [8], MetaMed [18], PT-MAP [48], Baseline [47], NegMargin [54], PFEMed [25].
Table 6: Comparison of our models and the SOTA methods. Values in the table are Accuracies (%) of the corresponding models on the Derm7pt dataset.
@@ 1 Shot 5 Shot
PCN Conv4 59.98±1.28 70.62±13
SCAN 61.42±1.49 72.58±1.28
Meta-DermDiagnosis Conv6 61.8 76.9
SCAN 62.80±1.34 76.65±1.21
NCA WRN-28-10 56.32±1.29 67.18±1.15
Baseline 59.43±1.34 74.28±1.14
S2M2_R 61.37±1.33 79.83±1.34
NegMargin 58.00±1.44 70.12±1.30
PT+NCM 60.92±1.68 74.33±1.48
PEMb E_NCM 60.40±1.72 72.63±1.48
EASY 61.02±1.67 75.98±1.41
SCAN 66.75±1.35 82.57±1.13
PFEMed 71.15 80.27
DTL-Base(Ours) DenseNet121 60.00±0.39 77.25±0.41
DTL-CutMix(Ours) 61.57±0.41 78.65±0.40
DTL-MixUp(Ours) 62.56±0.40 81.00±0.39
DTL-ResizeMix(Ours) 61.09±0.38 79.18±0.39
DTL-All-Augment(Ours) 60.88±0.40 78.30±0.40
Note: (1) The results of the SOTA models are taken from the SCAN [19]. (2) The complete references for the mentioned works in the table are as follows: PCN [14], Meta-DermDiagnosis [8], SCAN [19], NCA [53], Baseline [47], S2M2_R [46], NegMargin [54], PT+NCM [48], PEM bE-NCM [53], EASY [55], PFEMed [25].
In our research with the SD-198 dataset, we enhanced our Deep Transfer Learning (DTL) model by integrating various augmentation techniques, and compared these models with other state-of-the-art models in the field, particularly focusing on the SCAN model (Table 4). The SCAN model, which incorporates an unsupervised cluster branch, has been previously contrasted with the Meta-DermDiagnosis study. Our comparison also extended to various transfer learning-based Few-Shot Learning (FSL) methods; including the models like NCA [53], Baseline [47], S2M2_R [46], NegMargin [54], PT+NCM [48], PEMbE-NCM [53], and EASY [55]. Notably, the SCAN model reported better results than these approaches in tests on the SD-198 dataset. The EASY model, which achieved results comparable to the SCAN model, used an ensemble technique and augmented images with a random resize crop method. Our approach, however, is different; we employ the MobileNetV2 model, leveraging ImageNet pre-trained weights instead of the typically used WRN-28-10 model. Our augmentation strategy is more varied; we combined traditional techniques like Random Flip and Random Resize Crop with advanced methods such as CutMix, MixUp, and ResizeMix (as used by MetaMed model). This approach enabled our models to exceed the existing state-of-the-art results. Even though we used simple transfer learning based methodology, with the data augmentations our model could efficiently learn and transfer feature embeddings from base classes to novel classes.
Similarly, the PFEMed study compared its results on the ISIC2018 dataset with benchmarks set by studies like MetaMed [18], Meta-DermDiagnosis, PT-MAP [48], Baseline, and NegMargin. The PFEMed approach involved a comprehensive model with a dual-encoder structure combined with a Variational Autoencoder, distinguishing between general and specific features, and achieving state-of-the-art results. In contrast, the MetaMed study explored models employing augmentation techniques (such as CutOut, CutMix, and MixUp) during episodic training with Reptile model, and compared these to Meta-DermDiagnosis. In this dataset, observing the potential success of our DTL model trained with MobileNetV2, we applied similar augmentation techniques to the training data. The experiment results depict that our model performs comparable results with the state-of-the-art models; producing second best results in 3-Shot and 5-Shot cases. Moreover, we obtain best results when the number of samples is increased (i.e. 10-Shot case) (Table 5).
We also compared our models with leading models on the Derm7pt dataset. The unique features of this dataset played a key role in improving performance, especially the sub-cluster aspect in the SCAN model. We included findings from the PFEMed study in our comparison. By using the DTL-MixUp model, based on the DenseNet121 framework, we achieved results that align with existing benchmarks in the 2-Way 5-Shot scenario, where our model ranked second in accuracy. Similar to what we observed in our early experiments without data augmentations, in one-shot settings, the DTL model didn’t perform as well compared to other models (see Figure 5). It’s important to note that our method relies solely on transfer learning techniques and enhances the pretrained ImageNet model with data augmentations. These outcomes indicate that this fundamental approach remains effective in managing long-tail distributions, without the need for intricate training methods.

### v2

@@ 1 Shot 5 Shot 1 Shot 5 Shot
SCAN [29] Conv6 77.64±1.50 88.28±1.03 54.07±1.24 74.73±0.92
NCA [77] WRN-28-10 71.27±1.50 84.23±1.19 45.91±1.08 62.83±1.01
Baseline [71] 76.64±1.56 89.66±0.97 52.54±1.11 74.71±0.96
S2M2_R [70] 77.15±1.59 90.97±0.89 55.49±1.13 78.17±0.84
NegMargin [78] 77.98±1.45 90.65±0.92 56.04±1.14 77.75±0.87
PT+NCM [72] 78.86±1.47 90.90±0.93 56.91±1.11 78.12±0.88
PEMbE-NCM [77] 78.70±1.49 90.94±0.95 57.42±1.11 78.78±0.90
EASY [79] 79.44±1.51 91.43±0.96 57.77±1.12 79.53±0.89
SCAN [29] 81.21±1.46 92.08±0.85 58.75±1.14 81.43±0.77
DTL+Aug(Ours) ResNet50 75.04±0.38 94.11±0.62 55.93±0.43 86.55±0.71
DTL+Aug(Ours) ViT(Base) 81.76±0.65 92.10±0.45 64.00±0.37 85.10±0.59
DTL+Aug(Ours) DenseNet121 82.29±0.47 95.46±0.36 65.54±0.66 88.76±0.42
DTL+Aug(Ours) MobileNetV2 84.63±0.51 95.29±0.27 67.94±0.40 88.49±0.25
Note: The results of the SOTA models are taken from the SCAN [29].
5.2 Experimental Analysis
In order to evaluate our framework and assess the effectiveness of the methods, we conducted experiments using five few-shot configurations: 2-Way 1-Shot (2W1S), 2-Way 5-Shot (2W5S), 2-Way 10-Shot (2W10S), 5-Way 1-Shot (5W1S), and 5-Way 5-Shot (5W5S). These tests are conducted using three datasets: SD-198, Derm7pt, and ISIC2018. Model names in the results Table-2 follow a consistent notation where, for example, ResNet50(SimCLR) refers to a ResNet50 backbone that was first pretrained using the SimCLR self-supervised method and then fine-tuned on base classes for classification. The “+Aug” suffix indicates the application of batch-level data augmentation techniques such as MixUp, CutMix, or ResizeMix during fine-tuning. The consolidated results are presented in Tables 1 and 2. Table 1 summarizes the performances of backbone models without fine-tuning, while Table 2 presents the results of models fine-tuned under different training strategies, including FEL, FETL, DTL, DTLS, and baseline DL. Hence both the inherent capability of pretrained encoders and the improvements obtained through fine-tuning processes under few-shot learning settings is obtained.
Moreover, we executed additional experiments based on the parameters specified in Tables 3, 4, and 5. These experiments are designed to compare our model’s performance against benchmarks set in prior research, such as SCAN [29], MetaMed [28], and PFEMed [35], using the datasets referred to in Section [10, 11, 12]. In all these tables, the highest accuracy values are highlighted in bold, and the second highest results are underlined.
The SD-198 dataset contains 198 classes and consists only of clinical images. In contrast, the Derm7pt dataset has 18 classes with a mix of clinical and dermoscopic images. The ISIC2018 dataset, with its 7 classes, mainly features dermoscopic images. Although each of these three datasets shows a long-tail distribution, their unique features affect how models perform. In addition to dataset characteristics, model capacity in terms of parameter count and training computation cost (FLOPs) plays a significant role for few-shot classification performance. As detailed in Table 1, the tested models vary widely in size, ranging from lightweight architectures such as MobileNetV2 (3.5M parameters, 0.32G FLOPs) to heavyweight models like ViT-Base (86.6M parameters, 17.6G FLOPs). In our experiments, while compact models such as MobileNetV2 and DenseNet121 generally provided stable performance across few-shot setups due to their simplicity and lower risk of overfitting, larger models like ViT-Base demonstrated notably strong performance when appropriately fine-tuned on Derm7pt. These findings highlight that smaller capacity model alone is not a disadvantage in few-shot skin disease classification; rather, the effectiveness of large models depends on the availability of diverse and well-aligned data for adaptation.
Following the evaluation of Table 1, several backbone architectures, including MobileNetV2, DenseNet121, ViT-Base, and ResNet50, were selected for fine-tuning experiments due to their consistent performance across different datasets and shot configurations. These models demonstrated a favorable balance between model complexity and few-shot generalization capability. To further investigate the impact of contrastive self-supervised pretraining, we included ResNet50 models pretrained with methods such as SimCLR, SimSiam, and MoCo-v3 in our evaluations.
We have also included foundation models such as BioViL and MedCLIP in our evaluation by utilizing their vision encoders directly for few-shot classification. These models were originally pretrained on chest X-ray datasets using paired image–text data. When applied to our skin disease datasets, they performed below ImageNet-pretrained counterparts. This is likely due to a domain mismatch, as these models were trained on radiographic images with associated clinical text, whereas our datasets consist of dermoscopic and clinical skin images without accompanying text. Additionally, since we did not perform further fine-tuning due to lack of domain-specific textual data, their representations could not be adapted to the target domain, limiting their effectiveness in few-shot settings.
@@ 5.2 Experimental Analysis
First, across all datasets, FETL models consistently outperformed FEL models, particularly in lower-shot settings (e.g., 2W1S, 5W1S), highlighting the critical advantage of using pretrained ImageNet weights in few-shot scenarios. As the number of shots increased, transfer learning-based models, especially DTL and DTL, demonstrated significant gains over both FEL and FETL models. For instance, in the SD-198 dataset under the 5W5S setup, MobileNetV2 achieved 83.22% accuracy with FETL, while DTL with augmentation (DTL+Aug) reached 88.49%, clearly illustrating the impact of full fine-tuning and advanced data augmentation strategies. Similarly, DenseNet121 models showed a comparable trend: 84.02% with FETL and 88.76% with DTL+Aug. The behaviour of the models are similar in Derm7pt dataset. The reason behind this phenomenon lies in the adaptability of episodic learning, which performs better in scenarios with fewer shots due to the large number of classes. Additionally, clinical images inherently encapsulate various differences, making models trained in an episodic manner more inclined to tolerate these diversities. As the number of shots increases, models trained non-episodically become more successful in classifications.
When observing Derm7pt results, ViT-Base models adapted with LoRA fine-tuning exhibited outstanding performance, achieving over 93% accuracy in low-shot settings (2W5S and 5W5S). This suggests that large-capacity models, when fine-tuned properly on diverse datasets, can effectively overcome the challenges posed by few-shot classification.
The evaluation results presented in Table 2 demonstrate that among the contrastive pretraining strategies tested, SimCLR consistently yields the most favorable performance across all datasets when combined with augmentation and supervised fine-tuning (DTLS). For example, in the Derm7pt dataset, ResNet50(SimCLR)+Aug outperforms SimSiam and MoCo-v3 in nearly all configurations, particularly in 2W5S and 5W5S settings, where it achieves 80.58% and 63.26% accuracy, respectively. On ISIC2018, which has fewer classes and a more homogeneous modality, all contrastive methods perform comparably, but SimCLR still maintains a slight edge. These results suggest that SimCLR-based self-supervised pretraining provides a more transferable representation for downstream few-shot tasks, especially when followed by targeted fine-tuning with data augmentation. Overall, the choice of contrastive method can significantly affect performance, highlighting the importance of selecting an appropriate pretraining strategy for different dataset characteristics.
Table 4: Performance comparison of our models and SOTA methods on the ISIC2018 dataset, reported in accuracy (%).
Method Backbone 2 Way 3 Way
3 Shot 5 Shot 10 Shot 3 Shot 5 Shot 10 Shot
Meta-DermDiagnosis [9] Conv6 64.50 73.50 79.70 - - -
MetaMed - Transf. Learn. [28] Conv4 66.88 73.88 81.37 54.83 59.33 69.75
MetaMed - Normal Aug. [28] 72.75 75.62 81.37 54.83 59.33 69.75
MetaMed - CutOut [28] 70.37 77.62 81.87 55.50 65.41 69.75
MetaMed - MixUp [28] 75.37 78.25 84.25 58.50 61.25 71.00
MetaMed - CutMix [28] 73.25 76.87 80.62 58.66 61.50 66.50
PT-MAP [72] WRN 68.15 70.87 74.19 53.17 55.61 59.57
Baseline+ [71] 64.77 70.27 74.67 53.20 54.16 57.87
NegMargin [78] 71.33 72.67 75.17 60.69 57.58 63.04
Baseline [71] 68.77 71.03 76.97 56.80 59.20 65.22
PFEMed [35] 81.69 83.87 85.14 66.94 69.78 73.81
DTL+Aug (Ours) ResNet50 73.57 78.28 83.47 59.79 65.86 72.70
DTL+Aug (Ours) DenseNet121 62.42 65.06 67.85 45.74 48.66 52.34
DTL+Aug (Ours) ViT(Base) 62.85 65.64 67.98 47.22 49.55 53.25
DTL+Base (Ours) MobileNetV2 77.26 81.63 85.18 63.98 69.35 75.01
DTL+CutMix (Ours) 77.71 81.79 85.97 64.37 69.86 75.73
DTL+MixUp (Ours) 79.02 82.95 86.40 66.84 71.15 76.48
DTL+ResizeMix (Ours) 78.02 82.75 86.69 65.56 70.87 76.97
DTL+AllAug (Ours) 78.96 83.21 86.83 66.12 71.28 76.94
Note:The results of the SOTA models are taken from the PFEMed [35].
@@ 3 Shot 5 Shot 10 Shot 3 Shot 5 Shot 10 Shot
Since the ISIC2018 dataset comprises dermoscopic data, the discriminative power of domain-specific features becomes more crucial. Hence, transfer learning-based DTL and DL models outperform FETL and FEL models. For instance, for MobileNetV2, the accuracy rates are 58.49% for 2W1S with FETL and 64.83% with DTL, and 68.82% for 2W10S with FETL and 85.18% with DTL (Table 2). Episodic learning-based methods tend to remain superficial in training.
@@
Our proposed DTL approach exhibited promising performance on datasets containing clinical data, particularly SD-198. However, its performance was relatively lower on dermoscopic datasets such as ISIC2018 and, to some extent, Derm7pt. As illustrated in Table 2, DenseNet121-based models generally performed worse than MobileNetV2-based counterparts on ISIC2018, especially under low-shot conditions such as 2W1S and 3W1S. For example, under the 2W1S setup, DenseNet121 DTL achieved 67.51% accuracy, while MobileNetV2 DTL reached 69.5%. This gap can be attributed to DenseNet121’s tendency to overfit in low-data regimes. Moreover, ISIC2018’s structure (with only four base and three novel classes) offers limited variation for deeper architectures to generalize effectively. On Derm7pt, both MobileNetV2 and DenseNet121 showed more comparable results, with DenseNet121+Aug performing competitively in 5W5S. Nonetheless, across both datasets, neither model outperformed recent state-of-the-art methods such as SCAN or PFEMed. These results emphasize the need for improved techniques that specifically address the unique challenges of dermoscopic image classification, including fine-grained visual patterns and modality-specific artifacts. Notably, the ViT(Base)+LoRA+Aug model achieved the highest accuracy across several configurations on Derm7pt (e.g., 2W5S and 5W5S).
Our experiments further demonstrate that the transfer learning strategy in the FETL,DTL and DTLS methods effectively incorporates advances from natural image domains into skin disease classification tasks. Datasets include images with hair, ruler marks, varying body regions, and differences in skin tone. These artifacts may introduce instability in the models. However, leveraging data augmentation techniques like Flip and Resize, along with mix-based methods such as MixUp, CutMix, and ResizeMix, not only boosted model accuracy but also enhanced model robustness. A more detailed analysis and comprehensive discussion of these findings can be found in Section 6.
One limitation of our study is the absence of evaluation on independent clinical datasets, due to limited access beyond public benchmarks. While our models generalize well on standard datasets, future work will focus on external clinical validation to better assess real-world applicability.
5.3 Model Explainability
Figure 6: (a) Grad-CAM on SD-198, (b) Grad-CAM on Derm7pt, (c) Grad-CAM on ISIC2018, (d) t-SNE of 20 randomly selected SD-198 novel classes.
Figure 7: Prototype visualization with t-SNE on ISIC2018, Derm7Pt, and SD-198. Please refer to the digital version for better visibility.
Figure 8: Episode-level predictive uncertainty (mean entropy) versus episode accuracy for the SD-198 dataset with the MobileNetV2 model using Monte Carlo dropout.
To improve interpretability, we applied three common explainability techniques: prototype visualization with t-SNE [80], Grad-CAM [81], and predictive uncertainty estimation using Monte Carlo dropout [82]. All results were obtained using the DTL+Aug MobileNetV2 model across all datasets.
Grad-CAM maps (Figure 6), which are obtained from the last convolutional layer of the embedding extractor, show that the model generally attends to lesion regions in clinical datasets (SD-198 and Derm7pt), while attention in dermoscopic images (ISIC2018) is more scattered, which helps explain some misclassifications. Moreover, the t-SNE visualization of 20 randomly selected SD-198 classes in Figure 6(d) demonstrates that the base-trained model, despite never observing novel classes, is still able to form reasonably separable clusters in the embedding space.
In addition, we performed t-SNE visualization during testing using prototypical embeddings, particularly focusing on 2-way and 3-way episodes. Across all three datasets, distinct clustering is observed in the embedding space; however, clustering patterns are clearer in the clinical datasets (SD-198 and Derm7pt), while ISIC2018 remains less separable due to its higher visual similarity among classes (Figure 7).
Finally, the uncertainty analysis (Figure 8) revealed complementary insights at two levels. At the episode level, predictive entropy showed a negative correlation with accuracy (Pearson r=0.49r=-0.49 for 2-way 1-shot and r=0.64r=-0.64 for 2-way 5-shot). At the query level, misclassified samples tended to have higher entropy values, indicating that uncertainty is a useful signal for identifying unreliable predictions. Increasing the number of support samples from 1 to 5 reduced overall uncertainty and strengthened its correlation with accuracy, showing that predictions became more stable and entropy served as a more reliable error indicator. In addition to the overall decrease, entropy distributions became narrower with more support samples, indicating more stable predictions across episodes.
@@ 5.4 Comparison with Current State-of-the-Art
In our evaluation on the SD-198 dataset, we enhanced our Deep Transfer Learning (DTL) models using various data augmentation techniques and compared them against recent state-of-the-art methods, including SCAN, EASY, PT+NCM, and others (Table 3). Unlike the SCAN model, which uses a WRN-28-10 backbone with an unsupervised clustering branch, our models leverage lighter backbones such as MobileNetV2 and ViT, combined with ImageNet pretraining and multiple augmentation strategies (e.g., MixUp, CutMix, ResizeMix). Notably, our MobileNetV2-based DTL+AllAug model outperformed all prior methods in every configuration, achieving the highest F1-scores across all 2-Way and 5-Way shot settings. This indicates that even with a simpler architecture and traditional transfer learning, effective data augmentation can result in state-of-the-art performance.
On the ISIC2018 dataset, we compared our results with benchmarks established by PFEMed, MetaMed, PT-MAP, and others (Table 4). PFEMed, which incorporates a dual-encoder architecture with variational modeling, achieved strong results, particularly in low-shot 3-Way setups. However, our MobileNetV2 DTL+AllAug model achieved the highest accuracy in several 2-Way and 3-Way configurations, including 2W10S (86.83%) and 3W10S (76.94%). Furthermore, our ViT-Base-based model also showed competitive results, confirming that transformer backbones can generalize well in few-shot medical image settings when properly fine-tuned. These findings suggest that our approach can match or even surpass more complex meta-learning-based methods, particularly when sufficient training shots are available.
We further compared our models against state-of-the-art methods on the Derm7pt dataset. As shown in Table 5, our ViT(Base)+Aug DTL model achieved the highest accuracies in both 2-Way 1-Shot (86.35%) and 2-Way 5-Shot (93.68%) settings, outperforming strong baselines such as SCAN and PFEMed. In addition, MobileNetV2 and DenseNet121-based DTL models achieved competitive performance, especially when paired with augmentation techniques such as MixUp and CutMix. These results reinforce the notion that a well-designed transfer learning framework, which combines pretrained weights, architectural diversity, and batch-level augmentation, can effectively tackle the unique challenges posed by long-tailed dermatological datasets, without relying on meta-learning or external pretraining data.
Table 5: Performance of our models and SOTA methods on the Derm7pt dataset, reported in Accuracy (%).
@@ 1 Shot 5 Shot
PCN [24] Conv4 59.98±1.28 70.62±1.3
SCAN [29] 61.42±1.49 72.58±1.28
Meta-DermDiagnosis [9] Conv6 61.8 76.9
SCAN [29] 62.80±1.34 76.65±1.21
NCA [77] WRN-28-10 56.32±1.29 67.18±1.15
Baseline [71] 59.43±1.34 74.28±1.14
S2M2_R [70] 61.37±1.33 79.83±1.34
NegMargin [78] 58.00±1.44 70.12±1.30
PT+NCM [72] 60.92±1.68 74.33±1.48
PEMb E_NCM [77] 60.40±1.72 72.63±1.48
EASY [79] 61.02±1.67 75.98±1.41
SCAN [29] 66.75±1.35 82.57±1.13
PFEMed [35] 71.15 80.27
DTL+Aug (Ours) ResNet50 61.62±0.53 78.86±0.47
DTLS+Aug (Ours) ResNet50(SimCLR) 63.91±0.38 80.58±0.44
DTL+Aug (Ours) MobileNetV2 60.37±0.69 77.41±0.41
DTL+Aug (Ours) DenseNet121 62.56±0.40 81.00±0.39
DL (Ours) MoCo-v3(ViT) 88.40±0.55 92.18±0.63
DTL+Aug (Ours) ViT-Base 86.35±0.66 93.68±0.48
Note:The results of the SOTA models are taken from the SCAN [29].
6 Additional Analysis
We conducted additional analysis to examine how various augmentation strategies affect the performance of our DTL models on three benchmark datasets: SD-198, Derm7pt, and ISIC2018. We employed the following strategies: (1) Base (random horizontal and vertical flips at 45°), (2) CutMix, (3) MixUp, (4) ResizeMix, and (5) AllAug, which combines all the aforementioned methods on top of the Base augmentation. Detailed results for the SD-198, Derm7pt, and ISIC2018 datasets are presented in Tables 6.
We previously mentioned in Section 5.2 that artifacts in skin disease datasets compromise model stability, yet mix-based augmentation techniques effectively address these issues. Notably, the SD-198 dataset contains clinical images with varying artifacts (i.e. high scale variances within classes) due to different camera setups. Since ResizeMix resizes one image and overlays it onto another, it enhances the model’s resilience to such variations. Consequently, ResizeMix outperformed CutMix, MixUp, and AllAug by approximately 3% on the SD-198 dataset. In contrast, dermoscopy datasets, such as Derm7pt and ISIC2018, typically involve specialized imaging devices that capture deeper layers of the skin at uniform depths, minimizing distance-related variability and other artifacts. As a result, ResizeMix demonstrated similar performance to CutMix and MixUp on these two datasets.
Finally, we investigated how the architectural characteristics of MobileNetV2 and DenseNet121 influence the outcomes of these augmentations. Both backbones achieved comparable performance on the SD-198 and Derm7pt datasets. However, as discussed in the Discussion section, DenseNet121 suffered from overfitting on the ISIC2018 dataset and consequently failed to deliver robust results.
Table 6: Accuracy (%) comparison of models and augmentation techniques on the SD-198, Derm7Pt, and ISIC2018 datasets.
Dataset
Model
Method 2W-1S 2W-5S 5W-1S 5W-5S
SD-198
MobileNetV2
Base 82.42 94.40 64.06 86.67
Base+CutMix 83.00 94.77 65.48 87.63
Base+Mixup 82.29 94.48 64.56 87.22
Base+ResizeMix 84.63 95.29 67.94 88.49
Base+AllAug 83.31 95.02 65.99 88.04
SD-198
DenseNet121
Base 81.21 94.56 62.61 86.95
Base+CutMix 77.61 93.95 58.23 86.44
Base+Mixup 78.88 94.40 61.21 87.10
Base+ResizeMix 82.29 95.46 65.54 88.76
Base+AllAug 80.80 95.27 63.59 88.72
Derm7pt
MobileNetV2
Base 59.65 75.41 31.82 56.10
Base+CutMix 60.97 77.06 33.84 59.02
Base+Mixup 60.65 76.92 33.21 58.57
Base+ResizeMix 60.37 77.41 33.54 59.45
Base+AllAug 60.53 77.10 33.37 59.05
Derm7pt
DenseNet121
Base 60.00 77.25 32.04 57.68
Base+CutMix 61.57 78.65 35.46 60.58
Base+Mixup 62.56 81.00 34.75 61.91
Base+ResizeMix 61.09 79.18 33.89 61.22
Base+AllAug 60.88 78.30 33.19 59.95
Dataset
Model
Method 2W-1S 2W-5S 3W-1S 3W-5S
ISIC2018
MobileNetV2
Base 64.83 81.63 49.51 69.35
Base+CutMix 64.94 81.79 49.93 69.86
Base+Mixup 66.34 82.95 51.66 71.15
Base+ResizeMix 64.99 82.75 50.34 70,87
Base+AllAug 65.55 83.21 50.40 71.28
ISIC2018
DenseNet121
Base 55.55 64.69 38.63 46.47
Base+CutMix 54.45 60.78 38.63 44.32
Base+Mixup 55.20 64.42 39.72 48.98
Base+ResizeMix 55.90 63.57 39.85 48.27
Base+AllAug 56.67 65.06 39.81 48.66